Im a little lost on this word problem any help would be greatly appreciated.
A board 16ft long is cut into three pieces in such a way that the second piece is 1 ft longer then twice the first, and the third piece is 1ft shorter then the first. Find the length of each piece.
Thank You! Ana

Found 2 solutions by stanbon, tutorcecilia:
Answer by stanbon(48558)

A board 16ft long is cut into three pieces in such a way that the second piece is 1 ft longer then twice the first, and the third piece is 1ft shorter then the first. Find the length of each piece.
----------
Let the 1st piece have length "x" ft.
Then the 2nd piece has length "2x+1" ft.
And the 3rd piece has length "x-1" ft.
EQUATION:
1st + 2nd + 3rd = 16 ft.
x + 2x+1 + x-1 = 16
4x=16
x=4 ft.
1st piece=x=4ft
2nd piece= 2x+1=9 ft.
3rd piece+ x-1 = 3 ft.
Cheers,
Stan H.

Answer by tutorcecilia(2144)

If all the pieces were equal:
Piece #1 + Piece #2 + Piece #3 = 16 ft
and x+x+x=16
.
But the pieces are not equal because
Piece #1 = x
Piece #2=(2x+1)
Piece #3=(x-1)
So, let
x+(2x+1ft)+ (x-1ft)=16
x+2x+1+x-1=16
4x=16
x=4
.
Piece #1 = x=4
Piece #2=(2x+1)=(2*4+1)=9
Piece #3=(x-1)=(4-1)=3
So,
4+9+3=16
.
Check by plugging (x=4) back into the original equation.
x+(2x+1ft)+ (x-1ft)=16
(4)+(2(4)+1)+ ((4)-1)=16
16=16 [Checks out]